OpenStudy (anonymous):

whats sin 2/sin equivalent too.......1)cos, 2)2, 3)2sin,4) 2cos

OpenStudy (anonymous):

the expression sin^2 x + cos^2 x / cos x is equivalent to 1) sin x 2) cos x 3) sec x 4) csc x 4) if the probability that an event will occus is x/y , what

OpenStudy (anonymous):

3) the expression sin^2 x + cos^2 x / cos x is equivalent to 1) sin x 2) cos x 3) sec x 4) csc x sin^2 + cos^2 = 1 (a basic identity. Do you need proof of it?) sin^2 x + cos^2 x / cos x = 1/cos = secant(x) 3) sec(x) ---------- 4) if the probability that an event will occus is x/y , what is the probability that it will not occur? 1) y/x 2) -x/y 3) y - x / y 4) 1 - x / y All probabilities total 100%, or 1. So the probability it will not occur = 1 - x/y 4) 1 - (x/y)

OpenStudy (anonymous):

hope i helped mark the best response to show ur thanks